The product of two consecutive even integers is equal to 48. Translate into an equation and solve for factors.
Answer by Cromlix(4381)

Consecutive even integers = n, n + 2
n(n+2) = 48
n^2 + 2n - 48 = 0
(n + 8)(n - 6) = 0
n + 8 = 0
n = -8 (no answer n<0)
n - 6 = 0
n = 6
Therefore consecutive even numbers are:-
6 and 8
Hope this helps.
:-)